OpenStudy (anonymous):

how would you solve (squareroot363) (squareroot300)

OpenStudy (anonymous):

Well, we know couple rules: \(\sqrt{a^2b}=a\sqrt{b}\) \(\sqrt{a}\sqrt{b}=\sqrt{ab}\) So let us take a look, we have this expression: \[\sqrt{363}\sqrt{300}\] I will simplify this to \[a\sqrt{300}\] Let us take a closer look at \(\sqrt{363}\), can we factor anything out of it to break it up into \(\sqrt{a\times\frac{363}{a}}\) ?

OpenStudy (anonymous):

Think: (lets start at 2, not 1, lol) Can 2 go into it? can 3 go into it? can 4 go into it? Tell me what you see :). When you get to a whole number which divides into 363 and leaves a whole number, let me know :)

OpenStudy (anonymous):

three goes into it

OpenStudy (anonymous):

Try it out! \[\frac{363}{2}=181.5\] That's not a whole number. Try \(\frac{363}{3}\) Is that a whole number?

OpenStudy (anonymous):

EXACTLY! So we can simplify \(\sqrt{363}\) to \(\sqrt{3\times\frac{363}{3}}=\sqrt{3\times121}=\sqrt{3}\sqrt{121}\) Notice what happens when you punch in \(\sqrt{121}\) Into a calculator

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Right. So we can establish: \[\sqrt{363}=\sqrt{3}\sqrt{121}=11\sqrt{3}\] Make sense?

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

Try doing the same thing to \(\sqrt{100}\) You cool?

OpenStudy (anonymous):

tis 300

OpenStudy (anonymous):

Ehh what is?

OpenStudy (anonymous):

\[\sqrt{300}\sqrt{363}\neq300\]

OpenStudy (raden):

sqrt(300) = sqrt(100 * 3) = sqrt(100) * sqrt(3) = 10 sqrt(3)

OpenStudy (raden):

so, (squareroot363) (squareroot300) = 11 sqrt(3) * 10sqrt3) = 110 * sqrt(3) * sqrt(3) = 110 * 3 = 330

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Haha @RadEn ...Was trying not to throw the answer away but w/e lol

OpenStudy (raden):

hehe, just smack down this problem :)

OpenStudy (anonymous):

I would say ;)